Following information relates to Process A account 10000 Units introduced at Rs 20000 Material consumed Rs 5000 Manufacturing expenses Rs 2000 Direct Labour charges Rs 6000 Units transferred to next process B- 9500 It is bound that wastage in this process is 3% There is no realizable value of scrap Prepare the process account reflecting the value at which output will be transferred to the next process.


1
Expert's answer
2021-08-29T16:52:47-0400

Rs 20000 Material consumed

Rs 5000 Manufacturing expenses

Rs 2000 Direct Labour charges

Rs 6000 Units transferred to next process B- 9500


"5000+2000+6000+9500+0.03\times(5000+2000+6000+9500)=22500+675=23175"